OpenStudy (anonymous):

A store is charged $5.50 each for calculators and a delivery charge of $25 for the order. If the store sells the calculators for $8 each, how many calculators must be ordered and sold to produce a profit of atleast $80 dollars? (inequality word problem) (i need help making an equation, pleasee and thankyou!)

OpenStudy (anonymous):

Let x represent the number of calculators 5.50x + 25 = n (1) 8x = n + 80 (2) Substitute (1) into (2) and solve for x I guess that would be your answer.

OpenStudy (anonymous):

THANK YOU SO MUCH <3

OpenStudy (anonymous):

ahah make sure you do it to see if it's right :P Aha no problem